OpenStudy (anonymous):

The circumference of a circle with radius 2.3 cm is ____ cm to the nearest whole number.

OpenStudy (shamim):

formula for circumference of a circle is\[2 \pi r\]

OpenStudy (anonymous):

I got that but how do you plug it in

OpenStudy (shamim):

r=2.3 cm

OpenStudy (shamim):

\[2 \times \pi \times 2.3\]

OpenStudy (shamim):

now use calculator

OpenStudy (anonymous):

i got 14.45

OpenStudy (shamim):

14.451 cm

OpenStudy (shamim):

wts the wrong

OpenStudy (anonymous):

aw ok i get it

OpenStudy (shamim):

ok

OpenStudy (anonymous):

thank you

OpenStudy (shamim):

welcome

OpenStudy (shamim):

medal?

OpenStudy (anonymous):

yea